OpenStudy (anonymous):

Travis deposited $100 into a savings account that earns 6% compounded monthly. He plans on keeping the money in the account for 6 months and depositing an additional $100 each month. at the end of 6 months, how much interest will Travis have earned?

OpenStudy (tkhunny):

What is the monthly interest rate?

OpenStudy (anonymous):

A) 17.00 B) 10.50 C) 10.25 D)12.75 these are the possible answers.

OpenStudy (anonymous):

It was a closed question that was not answered.

OpenStudy (tkhunny):

It is of no value to know those. Answer the question.

OpenStudy (anonymous):

Well if these are the choices then you will know if you are wrong if you do it incorrectly.

OpenStudy (anonymous):

And this is asking about monthly compounding interest not simple interest.

OpenStudy (tkhunny):

Ha! I prefer to do it correctly. Now, answer my question. What is the MONTHLY interest rate?

OpenStudy (anonymous):

Principal starting is 100.00 Interest Rate is 6% or 12/6 rate is 0.0050 1st month interest is .50, Second month Principal grows to 200 and interest is 1.00 third month principal is 300 and interest is 1.50, 4th Months principal is 400 and interest is 2.00, 5th months principal is 5.00 and interest is 2.50, 6th months principal is 600 and interest is 3.00 add up the interest earned for each month So the answer is $10.50 interest earned over 6 months. ANSWER B above.

OpenStudy (anonymous):

tkhunny, you did not even attempt the problem so how did you do it? Having answers to choose from helps when working out possible answers.

OpenStudy (anonymous):

And I do not have my calculator either.

OpenStudy (tkhunny):

Having "answers" usually wastes time and leads to confusion. Please just answer the question. What is the MONHTLY interest rate? We know the annual rate is 6%. We also know compounding will be monthly. You will need a calculator of some sort. You can find one online or your phone may be sufficient.

OpenStudy (anonymous):

It was not my question the monthly interest rate is 6% divided by 12 as there ar 12 months in a year and it changes according to the amount in the account each month. Don;t need a calculator for that.

OpenStudy (anonymous):

And on my state test coming up we are NOT allowed to use a calculator on the Math test.

OpenStudy (anonymous):

Monthly interest rate is 0.0050

OpenStudy (tkhunny):

Ah! So the monthly interest rate is i = 0.06/12 = 0.005 It is very unlikely that you will be able to solve this sort of problem in a reasonable time without a calculator. Will there be an interest rate sheet or anything?

OpenStudy (anonymous):

Have no idea but the stakes are high and I go to a class that teaches mental math and Chinese abacus on Saturdays and it is much quicker than calculator.

OpenStudy (anonymous):

Look at utube contest between calculator and abacus it is cool.

OpenStudy (tkhunny):

Very good. As long as we have something that will do the trick. Well, then let's just walk through six months. Time 0 Month: Time of 1st Deposit: $100.00 Time 1 Month: $100.00(1+i) + $100.00 = $100.50 + $100.00 = $200.50 Time 2 Months: $200.50(1+i) + $100.00 = $201.50 + $100.00 = $301.50 <== Just introduced rounding error. Shall we continue, or would you prefer the Geometric Series method? We can avoid rounding error.

OpenStudy (anonymous):

I am in 6th Grade even though it is spring break I have to go to bed. Bye.

OpenStudy (tkhunny):

Fair enough. Past my bedtime, too.